OpenStudy (anonymous):

Convert into vertex form: y = 2(x - 3)(x + 4)

OpenStudy (anonymous):

y = 2(x^2 -3x + 4x -12) y = 2x^2 +2x -24 y = 2(x^2+x - 12) y = 2(x^2+x+0.25 -12.25) y = 2(x^2+x+0.25) - 24.5 y = 2(x+0.5)^2 - 24.5